OpenStudy (anonymous):

What is the solution set for |-2x+8|=12 A. x=-2 and x=10 B. x=-2 and x=-10 C. x=2 and x=10 D. x=2 and x=-10

OpenStudy (anonymous):

@raebaby420

OpenStudy (anonymous):

-2x+8=12 and -2x+8=-12 -2x=4 and -2x=-20 x=-2 and x=10

OpenStudy (anonymous):

@raebaby420

OpenStudy (anonymous):

That was a great example!! Thank you!!

OpenStudy (anonymous):

u are welcome